# Revealing hidden structures and symmetries in nonequilibrium transport

### 25 Jan 2024 v1

@@ Abstract
Recent results have shown how to partition the parameter space of Markov systems into dynamical equivalence classes. These equivalence classes structure transport properties in such a way that makes, among other features, their responses fully symmetric. In this note, I illustrate this approach on two representative systems. First, I derive analytical expressions for the equivalence classes of a disordered ring model. Second, I verify on a model of ion transport that, within an equivalence class, the response of coupled currents is symmetric both near and far from equilibrium.
Context and objectives
@@
However, despite their relevance, these core nonequilibrium concepts are not enough to build a well-defined nonequilibrium theory. What do I mean by that?
Take a nonequilibrium system of your choice. It is, by definition, out of equilibrium - but with respect to what? There is a continuum of equilibrium states that could serve as a reference point. And without a clearly defined equilibrium state, we cannot meaningfully characterize the nonequilibrium response of the system as there is no unique way to vary the set of parameters.
Choosing an arbitrary reference equilibrium state doesn’t solve the issue either. From a given equilibrium state, there is an infinite number of ways you can drive the system out of equilibrium. While in the linear response regime it doesn’t really matter since all response coefficients are identical regardless of the nonequilibrium paths you choose, this is no longer true in the nonlinear regime [1].
@@
  • Well-defined, i.e. there is a unique way to vary a system’s parameters to generate all possible nonequilibrium conditions
  • Intrinsic, i.e. the association emerges from the dynamics of the system itself.
To structure the space of possible dynamics while satisfying these requirements, I previously introduced the concept of dynamical equivalence classes [2, 3]. In addition to offering a well-defined nonequilibrium theory, they also bring extra benefits. For example, the current fluctuations can be obtained from an equilibrium dynamics and their nonlinear responses takes an especially simple form [4]. These results go beyond the traditional current fluctuation theorem [4, 2].
This new framework thus offers a new lens to understand nonequilibrium systems. Yet, it remains difficult to form a physical intuition around the equivalence classes themselves. This motivated my recent analysis of a molecular motor toy model, for which I derived exact expressions for its equivalence classes and transport properties [5].
In this note, my objective is to provide additional examples of how this framework plays out for specific systems. I first provide an analytical expression for the equivalence classes of a disordered ring model. Next, using a model of ion transport in membranes, I illustrate the symmetries of the response coefficients when multiple currents are present. I hope these examples will help to build physical intuition and kickstart the broader application of this framework, e.g. to the design and control of nonequilibrium systems.
I Transport on a Disordered Ring
@@
differs from zero and the system is out of equilibrium. From this expression, it is already clear that there is no unique way to define a nonequilibrium response. There is an infinite numbers of paths in $\Sigma$ with the same affinities (and starting from the same equilibrium state) but with different currents [5].
What we can do instead is to partition the space $\Sigma$ into dynamical equivalent classes [2]. Two dynamics $P$ and $Q$ belong to the same class if there exists a scalar $\rho$ such that:
$P\sim Q\quad{\rm if}\quad P\circ P^{T}=\rho\,\,Q\circ Q^{T}$ (10)
@@
We can use relation (10) to generate the equivalence classes of the disordered ring (7). Let’s consider an equilibrium dynamics $Q=\bar{Q}$ with transition probabilities $\bar{q}_{i}$ satisfying (8). Any dynamics $P$ in $[Q]$ then satisfies the system of $N$ polynomial equations
@@
and where $\rho$ varies between $[0,1]$ and the coefficients $a_{i}$ satisfy periodic boundary conditions, i.e. $a_{N+1}=a_{1}$ and $a_{0}=a_{N}$. In this parametrization, equilibrium occurs when $\rho=1$ or equivalently $\Delta=0$ [6]. For $\rho<1$ ($\Delta\neq 0$), both solutions $p^{\pm}_{i}$ belong to $[\bar{P}]$. They are time-reversed of each other, with opposite affinities and currents.
@@
We thus see that dynamical equivalence classes have a clear physical interpretation. For the disordered ring (7), each equivalence class determines a curve in $\Sigma$ that crosses the equilibrium hypersurface only once and whose affinities span the entire range [,][$\infty$,$\infty$][ - ∞ , ∞ ] as the curves tends to $p_{i}=0$ or $p_{i}=1$ for all $i$ (Figure 1).
As a result, within an equivalence class the nonequilibrium behavior is well-defined: there is a unique equilibrium state [10], all possible affinities are spanned, and there is a unique way to vary the parameters while driving the system out of equilibrium. Moreover, for each dynamics in $[\bar{Q}]$ its time-reversed one also belongs to $[\bar{Q}]$ with opposite affinities and currents. These constitute general properties of dynamical equivalence classes [4, 5].
Now that we understand the structure of equivalence classes and their physical interpretation, we can use them to study the nonequilibrium behavior of the disordered ring system.
@@
The higher-order cumulants $K^{(n)}(A)$ can then be directly obtained from the mean current as
$K^{(n)}(A)=\frac{{\rm d}J^{n}(A)}{{\rm d}A}\,.$ (15)
Note that, within an equivalence class, this relation is valid outside the linear regime and at all orders $n$. In contrast, the traditional fluctuation relations take a more complex forms [4].
@@
The solution (14) of the disordered ring offers a window into the nonequilibrium dynamics of systems with one independent current. For systems for more than one currents, additional symmetries emerge within an equivalence class [4]. We illustrate these additional symmetries in the next section.
II Transport of ions through a membrane
Let’s now turn to a model of ions transport as an archetype for systems with two independent currents. Following Hill [8], consider a cell surrounded by a membrane that separates the cell’s interior (In) from its environment (Out). A complex E, which can exist in two conformations E and E*, has binding sites for ions L and M. These sites are accessible to inside molecules in configuration E only, and to outside molecules in configuration E* only. L can be bound only when M is already bound on its site (Figure 3a).
@@ II Transport of ions through a membrane
The system has $N=6$ states connected by the kinetic diagram depicted in Figure 3b. The parameter space $\Sigma$ is $8$-dimensional while the set of equilibrium dynamics form an $6$-dimensional manifold in $\Sigma$. Each equivalence class corresponds to a $2$-dimensional manifold [13, 14].
Transport is determined by two independent currents, here measured along cycles $a$ and $b$ (Figure 3c). In Figure (4), we depict the currents $J_{a}(\boldsymbol{A})$ and $J_{b}(\boldsymbol{A})$ as a function of the affinities $A_{a}$ and $A_{b}$ in a given equivalence class. As expected, the currents are coupled, i.e. a change in $A_{a}$ will drive the current $J_{b}$ and vice versa. They display a nonlinear behavior and, beyond the anti-symmetry $\boldsymbol{J}(-\boldsymbol{A})=-\boldsymbol{J}(\boldsymbol{A})$ characteristic of equivalence classes, no other symmetry is apparent.
@@
Remarkably, despite the nonlinear behavior of the currents, their response is fully symmetric [5] both near and far from equilibrium:
@@
for all values of the affinities $\boldsymbol{A}$ in a given equivalence class (Figure 5). The symmetry (16) is also confirmed for all other equivalence classes (a Scilab code to generate the response coefficients for an arbitrary equivalence class is available upon request). Higher-order fluctuations can be obtained by further derivation of the currents and their responses are also symmetric [5].
@@ III Outlook
Properties like (15) and (16) confirm the importance of dynamical equivalence classes as a central concept in nonequilibrium thermodynamics. Where do these remarkable properties come from?
In the traditional way to look at nonequilibrium systems, affinities are varied along an arbitrary path in the parameter space (typically by keeping some parameters constant while varying the others). This makes every dynamics along the path unrelated to each other and with possibly very different dynamical properties. In contrast, within an equivalence class, all systems share common dynamical properties through relation (10). As a result, within an equivalence class the current generating functions of the different dynamics all have the exact same shapes, modulo a translation [2, 3]. All the fluctuations and reponse coefficients are thus directly related to each other, and equilibrium and nonequilibrium fluctuations are essentially identical.
Dynamical equivalence classes thus offer new insights by providing a well-defined structure and additional symmetries for nonequilibrium systems. Beyond their theoretical interests, they bring intriguing perspectives for the design and control of nonequilibrium systems.

### 19 Feb 2024 v2

@@ Abstract
Recent results have shown how to partition the space of Markov systems into dynamical equivalence classes. These equivalence classes structure transport properties in a way that makes, among other features, their responses fully symmetric. In this note, I illustrate this approach on two representative systems. First, I derive analytical expressions for the equivalence classes of a disordered ring model. Second, I verify on a model of ion transport that, within an equivalence class, the response of coupled currents is symmetric both near and far from equilibrium.
The Challenge: Nonequilibrium Response is ill-defined
@@
However, despite their relevance, these concepts are not enough to build a well-defined nonequilibrium theory. What do I mean by that?
The core issue is that specifying a system’s currents and affinities (and thus its entropy production) does not uniquely characterize its nonequilibrium state. Multiple combinations of parameters can lead to the same affinities and currents [1]. As a result, you can drive a system out of equilibrium in an infinite number of ways starting from a given equilibrium state. While in the linear response regime this ambiguity doesn’t really matter since all response coefficients are identical regardless of the path you choose, this is no longer true in the nonlinear regime [2].
Conversely, imagine now that you start with a system out of equilibrium. A continuum of equilibrium states could serve as reference point and, without a clearly identified equilibrium state and path to it, the nonequilibrium response of the system cannot be uniquely defined.
@@
  • Well-defined, i.e. the system’s parameters uniquely map all possible nonequilibrium conditions
  • Intrinsic, i.e. the connection emerges from the dynamical characteristics of the system itself.
To structure the space of possible dynamics while satisfying these requirements, I previously introduced the concept of dynamical equivalence classes [3, 4]. These ’hidden structures’ uniquely characterize the nonquilibrium conditions and parameters of the system, and describe a well-defined nonequilibrium theory. In addition, using these structures the current fluctuations can be obtained from an equilibrium dynamics and their nonlinear responses take an especially simple, symmetric form [5]. These properties go beyond the traditional current fluctuation theorem [5, 3], offering a new lens to understand nonequilibrium systems.
To build physical intuition around dynamical equivalence classes, I recently derived their structure and corresponding transport properties for a molecular motor toy model [6]. In this note, my objective is to provide additional examples of how this framework plays out in two specific cases. I first provide an analytical expression for the equivalence classes of a disordered ring model. Next, using a model of ion transport in membranes, I illustrate the symmetries of the response coefficients when multiple currents are present. I hope these examples will help to build physical intuition and kickstart the broader application of this framework, e.g. to the design and control of nonequilibrium systems.
I Hidden Structure: Transport on a Disordered Ring example
@@
differs from zero and the system is out of equilibrium. We see from this expression that the nonequilibrium response is not uniquely defined: An infinite number of paths exist in $\Sigma$ with the same affinities (and starting from the same equilibrium state) but with different currents [6].
To address this ambiguity, we can partition the space $\Sigma$ into dynamical equivalent classes as follows [3]. Two dynamics $P$ and $Q$ belong to the same class if there exists a scalar $\rho\geq 0$ such that:
$P\sim Q\quad{\rm if}\quad P\circ P^{T}=\rho\,\,Q\circ Q^{T}\,,$ (10)
@@
We can use relation (10) to generate the equivalence classes of the disordered ring (7). Consider an equilibrium dynamics $\bar{Q}$ with transition probabilities $\bar{q}_{i}$ satisfying (8). Any dynamics $P$ in $[\bar{Q}]$ then satisfies the system of $N$ polynomial equations
@@
and where $\rho$ varies between $[0,1]$ and the coefficients $a_{i}$ satisfy periodic boundary conditions, i.e. $a_{N+1}=a_{1}$ and $a_{0}=a_{N}$. In this parametrization, equilibrium occurs when $\rho=1$ or equivalently $\Delta=0$ [7]. For $\rho<1$ ($\Delta>0$), both solutions $p^{\pm}_{i}$ belong to $[\bar{P}]$. They are time-reversed of each other, with opposite affinities and currents.
@@
Dynamical equivalence classes thus have a clear physical interpretation. Each equivalence class determines a curve in $\Sigma$ (more generally a manifold) that crosses the equilibrium hypersurface only once and whose affinities span the entire range [,][$\infty$,$\infty$][ - ∞ , ∞ ] as the curves tends to $p_{i}=0$ or $p_{i}=1$ for all $i$ (Figure 1).
As a result, within an equivalence class the nonequilibrium behavior is well-defined: there is a unique way to vary the parameters while driving the system out of equilibrium. Moreover, for each dynamics in $[\bar{Q}]$ its time-reversed one also belongs to $[\bar{Q}]$ with opposite affinities and currents. These properties are valid for the equivalence classes of an arbitrary system [5, 6].
Now that we understand the structure of equivalence classes and their physical interpretation, let’s use them to study the nonequilibrium behavior of the disordered ring system.
@@
The higher-order cumulants $K^{(n)}(A)$ can then be directly obtained from the average current as
$\frac{1}{2^{n-1}}K^{(n)}(A)=\frac{{\rm d}J^{n}(A)}{{\rm d}A}\,.$ (15)
This relation holds outside the linear regime and at all orders $n$. In addition, knowledge of the current $J$ fully determines all the cumulants $K^{(n)}$. In contrast, the traditional fluctuation relations determine only half of the degrees of freedom and lead to more complex relationships across cumulants [5, 17].
@@
The solution (14) of the disordered ring provides insights into the nonequilibrium dynamics of systems with one independent current. For systems for more than one currents, the equivalence classes become higher-dimensional manifolds where additional symmetries emerge [5]. We illustrate these additional symmetries in the next section.
II Hidden symmetries: Ion Transport example
Let’s now turn to a model of ion transport as an archetype of systems with two independent currents. Following Hill [2], consider a cell surrounded by a membrane that separates the cell’s interior (In) from its environment (Out). A complex E, which can exist in two conformations E and E*, has binding sites for ions L and M. These sites are accessible to inside molecules in configuration E only, and to outside molecules in configuration E* only. L can be bound only when M is already bound on its site (Figure 3a).
@@ II Hidden symmetries: Ion Transport example
The system has $N=6$ states connected by the kinetic diagram depicted in Figure 3b. The parameter space $\Sigma$ is $8$-dimensional while the set of equilibrium dynamics forms an $6$-dimensional manifold in $\Sigma$.
Each equivalence class corresponds to a $2$-dimensional manifold [13, 14] related to the two independent affinties or currents (here measured along cycles $a$ and $b$, see Figure 3c). As expected, these currents are coupled, i.e. a change in $A_{a}$ will drive the current $J_{b}$ and vice versa (Figure 4). They display a nonlinear behavior and, beyond the anti-symmetry $\boldsymbol{J}(-\boldsymbol{A})=-\boldsymbol{J}(\boldsymbol{A})$ characteristic of equivalence classes, no other symmetry is apparent.
@@
Despite the nonlinear behavior of the currents, their response displays a ’hidden symmetry’ both near and far from equilibrium [6]:
@@
for all values of the affinities $\boldsymbol{A}$ in a given equivalence class (Figure 5). The symmetry (16) is confirmed for all other equivalence classes (a Scilab code to generate the response coefficients for an arbitrary equivalence class is available upon request). Systems with more than two independent currents display the same symmetries. Higher-order fluctuations can be obtained by further derivation of the currents and their responses are also symmetric [6].
@@ III Outlook
Properties such as (15) and (16) confirm the importance of dynamical equivalence classes as a central concept in nonequilibrium thermodynamics. Where do these remarkable properties come from?
In the traditional way of looking at nonequilibrium systems, affinities are varied along an arbitrary path in the parameter space (typically by keeping some parameters constant while varying the others). This makes each dynamics along the path unrelated to each other and with possibly very different kinetic properties. In contrast, within an equivalence class, all systems share similar kinetic properties through relation (10). Their current generating functions all have the same shapes, modulo a translation [3, 4]. All the fluctuations and response coefficients are thus directly related to each other, and equilibrium and nonequilibrium fluctuations are essentially identical.
Beyond their theoretical interests, dynamical equivalence classes bring intriguing perspectives for the design and control of nonequilibrium systems. For example, designing mesoscopic devices around the symmetry (16) could make them easier to control and/ or robust to perturbations.
More generally, suppose you operate a mesoscopic device designed to achieve a certain output (i.e., currents). The device’s parameters presumably reflect a trade-off around the desired transport properties and the admissible fluctuactions and dissipation. Imagine now that you need to achieve a different output while minimizing any other disruption to the device. How would you ensure the devive continues to operate as smoothly as possible?
One way to go about it would be to change the device’s outputs while also staying as close as possible to the starting dynamics to minimize unrelated deviations or disruptions. This is precisely what dynamical equivalence classes accomplish! If the device has starting dynamics $G$, its equivalence class is obtained as [12]
$\min_{P}\left[D(P|G)+\sum_{a}\lambda_{a}J_{a}[P]\right]\,.$ (17)
Here the $\lambda_{a}$s act as control parameters or Lagrange multipliers for the constrained currents. Meanwhile, the divergence $D(P|G)=\sum_{i,j}\pi_{i}P_{ij}\ln(P_{ij}/G_{ij})$ measures the statistical distance between $P$ and $G$ (in information theoric terms, $D$ is the additional cost in nats required to encode $P$ using measure $G$).
When solving the minimization problem, the Lagrange multipliers relate to the affinities as $\lambda_{a}=(A_{a}-\bar{A}_{a})/2$, where $\bar{A}_{a}$ are the affinities of $G$. In other words, the minimization happens across the space of dynamics $P$ with given affinities $A_{a}=2\lambda_{a}+\bar{A}_{a}$.
Putting everything together, Eq. (17) reveals that equivalence classes select the dynamics minimizing the statistical distance to $G$ given fixed currents (or, equivalently, fixed affinities). In other words, using equivalence classes we can control a device’s output while minimizing its distortions with respect to a target dynamics $G$. I believe this has practical applications and I hope these observations will stimulate further exploration of these concepts.